OpenStudy (anonymous):

Which of the following is equivalent to 5 2/5 ÷ 10? 1 23/27 1/2 27/50 5 1/25

OpenStudy (anonymous):

5 2/5 = 5.4 5.4/10 = 0.54 = 27/50

OpenStudy (anonymous):

Thanks